OpenStudy (aravindg):

More doubts:

OpenStudy (aravindg):

1. I read that if a liquid of radius R is split into n smaller droplets,all of same size and Surface tension T then radius of each droplets is r=R(n)^-1/3 How do we get that?

OpenStudy (aravindg):

@Mashy

OpenStudy (anonymous):

well.. lets take one example if you have a drop of radius 5cm, and its splits into two smaller drops.. of equal radii, and surface tension.. what do you think is the radius of each drop?

OpenStudy (aravindg):

R/4?

OpenStudy (anonymous):

how did you get that?

OpenStudy (aravindg):

4r=R R=r/4

OpenStudy (anonymous):

4r = R ? how did you ge tthat? :P

OpenStudy (aravindg):

oh wait ...

OpenStudy (anonymous):

think of how would you solve the problem when the drop splits.. what quantity becomes half?

OpenStudy (aravindg):

Area?

OpenStudy (anonymous):

nope.. area totally depends on the shape.. by changing shape i can change area.. easily!

OpenStudy (aravindg):

Ah final Volume :D

OpenStudy (anonymous):

lolz. exactly.. regardless of what you do.. if a drop occupies certain volume.. you split it.. or do whatever, its the volume that gets split accordingly so now what shoudl be the radius in my problem?

OpenStudy (aravindg):

R 2^(-1/3)

OpenStudy (anonymous):

did you directly use your formula.. or used volume consideration to arrive at that? :P

OpenStudy (vincent-lyon.fr):

Just solve: \(V = nv\) i.e. \(\dfrac{4}{3}\pi R^3=n\times \dfrac{4}{3}\pi r^3\)

OpenStudy (aravindg):

@Mashy I didnt use the formula. I got the answer by equating Volume as vince showed above. Thanks for the help :)